# Exact Modeling of Power Spectrum Multipole through Spherical Fourier-Bessel Basis

### v1

@@ Abstract
Current and upcoming large-scale structure (LSS) surveys such as DESI Aghamousa et al. (2016), Euclid Amendola et al. (2018), and SPHEREx Doré et al. (2014) will measure the galaxy density field over increasingly larger volume, offering improved constraining power on effects that manifest on large scales, such as the primordial non-Gaussianity (PNG) Komatsu and Spergel (2001); Dalal et al. (2008); Slosar et al. (2008); de Putter and Doré (2017); Rezaie et al. (2023) and general relativistic (GR) effects Yoo et al. (2009); Yoo (2010); Challinor and Lewis (2011); Bonvin and Durrer (2011); Jeong et al. (2012); Elkhashab et al. (2022). With wider angular coverage, the commonly used global plane-parallel approximation, which assumes the same line-of-sight (LOS) for all galaxy pairs, breaks down when the galaxy separation becomes large, and wide-angle (WA) effects become important. Additionally, the Newtonian approximation to redshift space distortion (RSD)– an effect that arise from our measuring the redshifts and not the distances of galaxies – also breaks down as GR effects become important on large scales.
@@
The next best thing to the global plane-parallel approximation is a local one in which each galaxy pair is allowed to have a different LOS. This is achieved by the Yamamoto estimator for measuring the Fourier-space power spectrum multipoles (PSM). An efficient implementation of this is to choose one of the galaxy to be the pair LOS, called the end-point LOS Bianchi et al. (2015); Scoccimarro (2015); Beutler et al. (2019). The signal picked up by this estimator at large scales includes WA effects that are usually modeled either perturbatively as an expansion in the galaxy pair separation Reimberg et al. (2016); Castorina and White (2018); Beutler et al. (2019); Benabou et al. (2024); Beutler and McDonald (2021); Noorikuhani and Scoccimarro (2023); Paul et al. (2023), or non-perturbatively via an exact calculation of the configuration-space correlation function Tansella et al. (2018); Castorina and Di Dio (2022).
@@
In this Letter, we propose an alternative non-perturbative method for calculating the PSM that employs a natural basis for the curved sky: the spherical Fourier-Bessel (SFB) basis, which was proposed in the 90s Binney and Quinn (1991); Lahav (1993) and has recently gained traction Leistedt et al. (2012); Yoo and Desjacques (2013); Nicola et al. (2014); Lanusse et al. (2015); Samushia (2019); Zhang et al. (2021); Grasshorn Gebhardt and Doré (2021); Khek et al. (2022); Grasshorn Gebhardt and Doré (2023). As the eigenfunctions of the Laplacian in the spherical coordinates, the SFB basis preserves the geometry of the curved sky and naturally models the WA effects. Due to the separation of radial and angular modes in spherical coordinates, one can incorporate Newtonian RSD and GR effects more easily for the SFB power spectrum Yoo and Desjacques (2013); Zhang et al. (2021) than for the PSM Castorina and Di Dio (2022).
@@
Despite these advantages, the SFB power spectrum is a less developed formalism compared to the PSM, especially at the mildly non-linear scales ($k\gtrsim 0.05\,h/{\rm Mpc}$). Its estimator also suffers from an increased computational cost due to the large number of modes, and the fact that one cannot employ fast Fourier-transforms as is often used in Cartesian-based estimators Grasshorn Gebhardt and Doré (2021, 2023).
@@
In this work, we build upon Ref. Castorina and White (2018) and fully develop a mapping from the SFB power spectrum to the PSM, which allows us to take advantage of the strengths in both statistics: on large scales we can exactly model WA and GR effects using the SFB power spectrum, before transforming the result into the PSM, to take advantage of its efficient Yamamoto estimator implementation and better developed small-scale modeling. This mapping represents a new way of non-perturbatively modeling WA effects in the PSM, which we validate using log-normal mocks. Moreover, for the first time we obtain an exact analytic covariance for the PSM, handling off-diagonal mode-couplings due to WA effects in the covariance, without resorting to approximations such as those made in Ref. Wadekar and Scoccimarro (2020). Though formulated in the context of galaxy surveys, our results will also have implications for future wide-field intensity mapping surveys Hall et al. (2013); Liu et al. (2016); Viljoen et al. (2021).
@@
Power Spectrum Multipoles–We first briefly review the SFB formalism. The canonical SFB basis is composed of eigenfunctions of the Laplacian in spherical coordinates, namely the spherical Bessel functions of first kind $j_{\ell}(kx)$ and the spherical harmonics $Y^{*}_{\ell m}(\hat{\mathbb{x}})$. The SFB decomposition of the galaxy density field $\delta(\mathbb{x})$ is then:
@@
where $\delta^{K}$ is the Dirac-delta. We use the upper indices $a$ and $b$ to indicate the radial modes and the lower index for the angular mode of the gSFB power spectrum. When $a=b=\ell_{1}$, it reduces to the canonical SFB power spectrum $C_{\ell_{1}}(k_{1},k_{2})$. Note that the SFB power spectra adopt two values of $k$, and we call $k_{1}=k_{2}$ diagonal and $k_{1}\neq k_{2}$ off-diagonal.
@@
where $\mathcal{L}_{L}(\hat{\mathbb{x}}\cdot\hat{\mathbf{k}})$ is the Legendre polynomial and $I_{22}\equiv\int_{\mathbb{x}}W(\mathbb{x})^{2}$ is the normalization. Using the plane-wave expansion and evaluating the $\hat{\mathbf{k}}$ integral (see full derivation in the Appendix), we find the relationship between the PSM and the gSFB power spectrum:
@@
The above equation was first derived in Ref. Castorina and White (2018). Compared to their expression, our Eq. 8 accounts for the window convolution and is expressed in terms of the gSFB power spectrum (Eq. 4) instead of the gSFB mode (Eq. 2). Going beyond Ref. Castorina and White (2018), we will highlight the theoretical importance of Eq. 8 and, for the first time, enable its use for PSM modeling by developing a method to compute the gSFB power spectrum.
@@
a sum over all the angular modes of the canonical SFB power spectrum. Note that only the diagonal components of the canonical SFB power spectrum are present. These are the only modes that would exist in a homogeneous and isotropic Universe where there were no redshift evolution and LOS-dependent effects like RSD Grasshorn Gebhardt and Doré (2021). This is exactly what the PS monopole is, an average over the orientation of the Fourier mode ${\bf k}$, and it erases the redshift evolution information by integrating over the redshift bin.
Given the above, we would then expect that the off-diagonal components $C_{b}^{W}(k,k^{\prime})$ would be (partially) brought back for the higher multipoles in $P_{L}$. Indeed, this is achieved through the generalized SFB $C_{b}^{ab,\rm W}(k,k)$, where by construction, the off-diagonal information in $C_{b}^{W}(k,k^{\prime})$ is folded into the extra upper radial indices $a$ and $b$, keeping a single $k$ value.
@@
and $\Delta_{\ell}(k,x(z))$ is the angular projection kernel that would contain effects such as RSD, GR Bonvin and Durrer (2011); Challinor and Lewis (2011); Di Dio et al. (2013) and Fingers-of-god (FoG) effects Khek et al. (2022). On linear scales where the FoG effect is small, and in the Newtonian limit, the angular projection kernel becomes
@@
Figure 1: The multipoles $P_{L}(k)$. The green lines show the theoretical results of PSM, and the blue points show the mean and the $1\sigma$ error of the measured multipoles from the 10,000 full-sky mocks for $z=0.2-0.5$. In the lower panels, we show the differences between the measurements and the theoretical results, normalized by the monopole that sets the simulation precision. The grey bands indicate $1\%$ accuracy. Our theoretical results and measurements agree well at large scales.
The integrals in Eq. 10 and 11 can be computed through the procedure introduced in Refs. Khek et al. (2022) and Grasshorn Gebhardt and Doré (2023) where the authors computed the canonical SFB power spectrum. The only change we make is allowing radial and angular modes to have different orders $\ell\neq a$ when computing the gSFB. With this we evaluate the radially-convolved gSFB power spectrum and obtain the PSM using Eq. 8 for an example of the full-sky window (taking $z$ between 0.2 and 0.5). We plot the theory computation for the first five multipoles in solid lines in Fig. 1.
@@
Validation with Full-sky Mocks– To validate Eq. 8 and showcase its use for modeling WA effects, we compare our theoretical results with measurements from full-sky simulations. In particular, we choose to use the set of log-normal mocks generated in Ref. Benabou et al. (2024). With the goal of studying a perturbative calculation of WA effects, the authors of Ref. Benabou et al. (2024) generated 10,000 full-sky mocks in linear theory with Newtonian RSD, spanning $z=0.2$ to $0.5$. They also provide a new implementation for the following Yamamoto estimator, whose ensemble average is expected to be Eq. 5:
@@
To compare with full-sky mocks under Newtonian RSD, the theoretical modeling of galaxy fluctuation $\delta(\mathbb{x})$ needs to include the canonical Newtonian RSD term, the angular kernel of which is given in Eq. 12. We further need to include an additional velocity contribution scaled with $v/x$, which is generated from the Jacobian associated with the change of coordinates caused by Newtonian RSD Szalay et al. (1998); Pápai and Szapudi (2008); Yoo and Seljak (2015); Raccanelli et al. (2010). This velocity term is usually ignored for mocks or previous surveys with small angular coverage Beutler et al. (2019); Beutler and McDonald (2021); Castorina et al. (2019), but it must be included in the full-sky limit Raccanelli et al. (2010); Benabou et al. (2024). The angular projection kernel corresponding to this velocity term is
@@
where $\alpha(x)\equiv\partial[\ln x^{2}R(x)]/\partial\ln x$, and it becomes $\alpha=2$ for a uniform radial window in the redshift bin Yoo and Seljak (2015); Castorina and White (2018).
The mocks in Ref. Benabou et al. (2024) are generated with $f=0.71$ and $b_{1}=1.455$, and there is no redshift evolution within the redshift bin for the mocks. All calculations assume a best-fit Planck 2018 cosmology Planck Collaboration VI (2020). The results for the mean of the multipole measurements from the 10,000 mocks with $L$ from 0 to 4 are shown in Fig. 1. We see that our theory predictions agree exceptionally well with the measured multipoles at the largest scales. The disagreements at the small scales ($k\gtrsim 0.06\,h/{\rm Mpc}$) are due to the voxel windows used in the simulations Benabou et al. (2024) which we do not include in our modeling.
Gaussian Covariance– We next express the Gaussian covariance of PSM in terms of the gSFB power spectrum. Compared to Ref. Wadekar and Scoccimarro (2020), which had the state-of-art results for analytical Gaussian covariance of the PSM, we will calculate the exact covariance without resorting to any plane-parallel or LOS approximation. The results in Ref. Wadekar and Scoccimarro (2020) are only applicable in the flat sky limit and on scales where the window function is subdominant, while our expression is exact and can be applied at the largest scales in particular for measuring PNG and GR effects. Using the Yamamoto estimator of Eq. 14, the Gaussian covariance reads Wadekar and Scoccimarro (2020):
@@
Using Wick contractions on the 4-point function and expressing $F_{L}(k)$ with Eq. 15, we can follow the same procedure as the PSM case. To simplify the final expression, we further assume the separability of the window function and obtain:
$\textbf{C}^{\textup{G}}_{L_{1}L_{2}}(k_{1},k_{2})=(4\pi)^{4}\frac{(2L_{1}+1)(2L_{2}+1)}{I_{22}^{2}}\sum_{a,b,c,d,\ell_{1},\ell_{2}}i^{-a-c+b+d}(2a+1)\begin{pmatrix}a&L_{1}&b\\ 0&0&0\end{pmatrix}^{2}\begin{pmatrix}c&L_{2}&d\\ 0&0&0\end{pmatrix}^{2}$ (22)
$\qquad\Bigg{[}(2c+1)S_{b\ell_{1}d\ell_{2}}+(-1)^{L_{2}}(2d+1)S_{b\ell_{1}c\ell_{2}}\Bigg{]}C_{\ell_{1}}^{ad,{\rm R}}(k_{1},k_{2})C_{\ell_{2}}^{bc,{\rm R}}(k_{1},k_{2})\,,$ (23)
@@
Under the same set-up as our log-normal mocks with full-sky window, we plot an example of the analytical Gaussian covariance matrix using Eqs. 10 and 26 in Fig. 2, which shows prominent off-diagonal components with $k\lessapprox 0.02h/{\rm Mpc}$.

### v2

@@ I Introduction
Current and upcoming large-scale structure (LSS) surveys such as DESI Aghamousa et al. (2016), Euclid Amendola et al. (2018), and SPHEREx Doré et al. (2014) will measure the galaxy density field over increasingly larger volume, offering improved constraining power on effects that manifest on large scales, such as the primordial non-Gaussianity (PNG) Komatsu and Spergel (2001); Dalal et al. (2008); Slosar et al. (2008); de Putter and Doré (2017); Rezaie et al. (2024) and general relativistic (GR) effects Yoo et al. (2009); Yoo (2010); Challinor and Lewis (2011); Bonvin and Durrer (2011); Jeong et al. (2012); Elkhashab et al. (2022). With wider angular coverage, the commonly used global plane-parallel approximation, which assumes the same line-of-sight (LOS) for all galaxy pairs, breaks down when the galaxy separation becomes large, and wide-angle (WA) effects become important. Additionally, the Newtonian approximation to redshift space distortion (RSD)– an effect that arises from our measuring the redshifts and not the distances of galaxies – also breaks down as GR effects become important on large scales.
@@
The next best thing to the global plane-parallel approximation is a local one in which each galaxy pair is allowed to have a different LOS. This is achieved by the Yamamoto estimator for measuring the Fourier-space power spectrum multipoles (PSM). An efficient implementation of this is to choose one of the galaxy to be the pair LOS, called the end-point LOS Bianchi et al. (2015); Scoccimarro (2015); Beutler et al. (2019). The signal picked up by this estimator at large scales includes WA effects that are usually modeled either perturbatively as an expansion in the galaxy pair separation Reimberg et al. (2016); Castorina and White (2018); Beutler et al. (2019); Benabou et al. (2024); Beutler and McDonald (2021); Noorikuhani and Scoccimarro (2023); Paul et al. (2023), or non-perturbatively via an exact calculation of the configuration-space correlation function Tansella et al. (2018); Castorina and Di Dio (2022). In a companion paper, Ref. Benabou et al. (2024) demonstrates that the perturbative method for WA effects breaks down at large scales, which can bias constraints on PNG and motivate further study of non-perturbative WA modeling.
@@
In this work, we propose an alternative non-perturbative method for calculating the PSM that employs a natural basis for the curved sky: the spherical Fourier-Bessel (SFB) basis, which was proposed in the 90s Binney and Quinn (1991); Lahav (1993) and has recently gained traction Leistedt et al. (2012); Yoo and Desjacques (2013); Nicola et al. (2014); Lanusse et al. (2015); Samushia (2019); Wang et al. (2020); Zhang et al. (2021); Grasshorn Gebhardt and Doré (2021); Khek et al. (2024); Grasshorn Gebhardt and Doré (2024). As the eigenfunctions of the Laplacian in the spherical coordinates, the SFB basis preserves the geometry of the curved sky and naturally models the WA effects. Due to the separation of radial and angular modes in spherical coordinates, one can incorporate Newtonian RSD and GR effects more easily for the SFB power spectrum Yoo and Desjacques (2013); Zhang et al. (2021) than for the PSM Castorina and Di Dio (2022).
@@
Despite these advantages, the SFB power spectrum is a less developed formalism compared to the PSM, especially at the mildly non-linear scales ($k\gtrsim 0.05\,h/{\rm Mpc}$). Its estimator also suffers from an increased computational cost due to the large number of modes, and the fact that one cannot employ fast Fourier-transforms as is often used for PSM Grasshorn Gebhardt and Doré (2021, 2024).
@@
In this work, we build upon Ref. Castorina and White (2018) and fully develop a mapping from the SFB power spectrum to the PSM in Sec. II. This mapping allows us to take advantage of the strengths in both statistics: on large scales we can exactly model WA and GR effects using the SFB power spectrum, before transforming the result into the PSM, to take advantage of its efficient Yamamoto estimator implementation and better developed small-scale modeling. This mapping represents a new way of non-perturbatively modeling WA effects in the PSM, and in Sec. III, we validate this new approach to WA modeling using the estimator and log-normal mocks described in a companion paper, Ref. Benabou et al. (2024). Moreover, for the first time we obtain an exact analytic covariance for the PSM (Sec. IV), handling off-diagonal mode-couplings due to WA effects in the covariance, without resorting to approximations such as those made in Ref. Wadekar and Scoccimarro (2020). Though formulated in the context of galaxy surveys, our results will also have implications for future wide-field intensity mapping surveys Hall et al. (2013); Liu et al. (2016); Square Kilometre Array Cosmology Science Working Group et al. (2020); Viljoen et al. (2021).
II Power Spectrum Multipoles
@@
We first briefly review the SFB formalism. The canonical SFB basis is composed of eigenfunctions of the Laplacian in spherical coordinates, namely the spherical Bessel functions of first kind $j_{\ell}(kx)$ and the spherical harmonics $Y^{*}_{\ell m}(\hat{\mathbb{x}})$. The SFB decomposition of the galaxy density field $\delta(\mathbb{x})$ is then:
@@
where $\delta^{K}$ is the Dirac-delta. We use the upper indices to indicate the radial modes and the lower index for the angular mode of the gSFB power spectrum. When $a=b=\ell_{1}$, it reduces to the canonical SFB power spectrum $C_{\ell_{1}}(k_{1},k_{2})$. Note that the SFB power spectra adopt two values of $k$, and we call $k_{1}=k_{2}$ diagonal and $k_{1}\neq k_{2}$ off-diagonal.
@@
where $\mathcal{L}_{L}(\hat{\mathbb{x}}\cdot\hat{\mathbf{k}})$ is the Legendre polynomial and $I_{22}\equiv\int_{\mathbb{x}}W(\mathbb{x})^{2}$ is the normalization. Using the plane-wave expansion and evaluating the $\hat{\mathbf{k}}$ integral (see full derivation in the Appendix. A), we find the relationship between the PSM and the gSFB power spectrum:
@@
The above equation was first derived in Ref. Castorina and White (2018). Compared to their expression, our Eq. 6 accounts for the window convolution and is expressed in terms of the gSFB power spectrum (Eq. 4) instead of the gSFB mode (Eq. 2). Going beyond Ref. Castorina and White (2018), we will highlight the theoretical importance of Eq. 6 and, for the first time, enable its use for theoretical modeling of PSM by developing a method to compute the gSFB power spectrum.
@@
a sum over all the angular modes of the canonical SFB power spectrum. Note that only the diagonal components of the canonical SFB power spectrum are present. These are the only modes that would exist in a homogeneous and isotropic Universe where there are no redshift evolution and LOS-dependent effects like RSD Grasshorn Gebhardt and Doré (2021). This is exactly what the PS monopole is, an average over the orientation of the Fourier mode ${\bf k}$, and it erases the redshift evolution information by integrating over the redshift bin.
Given the above, we would then expect that the off-diagonal components $C_{b}^{\rm W}(k,k^{\prime})$ would be (partially) brought back for the higher multipoles in $P_{L}$. Indeed, this is achieved through the generalized SFB $C_{b}^{ab,\rm W}(k,k)$, where by construction, the off-diagonal information in $C_{b}^{\rm W}(k,k^{\prime})$ is folded into the extra upper radial indices $a$ and $b$, keeping a single $k$ value.
@@
and $\Delta_{\ell}(k,x(z))$ is the angular kernel that would contain effects such as RSD, GR Bonvin and Durrer (2011); Challinor and Lewis (2011); Di Dio et al. (2013) and Fingers-of-god (FoG) effects Khek et al. (2024). On linear scales where the FoG effect is small, and in the Newtonian limit, the angular kernel becomes
@@
Figure 1: The theoretical and measured multipoles $P_{L}(k)$. The green lines show the theoretical results of PSM, and the blue points show the mean and the $1\sigma$ error of the measured multipoles from the 10,000 full-sky mocks for $z=0.2-0.5$ based on the Yamamoto estimator. The $1\sigma$ errors shown here represent the error on the mean instead of the error for the measured spectrum from a single realization, and the errors are obtained by calculating the standard deviation of the 10,000 measured spectra. In the lower panels, we show the differences between the measurements and the theoretical results, normalized by the monopole that sets the simulation precision. The grey bands indicate $1\%$ accuracy. Our theoretical results and measurements agree well at large scales.
The integrals in Eq. 8 and 9 can be computed through the procedure introduced in Refs. Khek et al. (2024) and Grasshorn Gebhardt and Doré (2024) where the authors computed the canonical SFB power spectrum. The only change we make is allowing radial and angular modes to have different orders $\ell\neq a$ when computing the gSFB. With this we evaluate the radially-convolved gSFB power spectrum and obtain the PSM using Eq. 6 for an example of the full-sky window (taking $z$ between 0.2 and 0.5). We plot the theory computation for the first five multipoles in solid lines in Fig. 1.
@@
III Validation with Full-sky Mocks
To validate Eq. 6 and showcase its use for modeling WA effects, we compare our theoretical results with measurements from full-sky simulations. In particular, we choose to use the set of log-normal mocks generated in Ref. Benabou et al. (2024). With the goal of studying a perturbative calculation of WA effects, the authors of Ref. Benabou et al. (2024) generated 10,000 full-sky mocks in linear theory with Newtonian RSD, spanning $z=0.2$ to $0.5$. They also provide a new implementation for the following Yamamoto estimator, whose ensemble average is expected to be Eq. 5:
@@ III Validation with Full-sky Mocks
To compare with full-sky mocks under Newtonian RSD, the theoretical modeling of galaxy fluctuation $\delta(\mathbb{x})$ needs to include the canonical Newtonian RSD term, the angular kernel of which is given in Eq. 10. We need to include an additional velocity contribution scaled with $v/x$, which is generated from the Jacobian associated with the change of coordinates caused by Newtonian RSD Szalay et al. (1998); Pápai and Szapudi (2008); Yoo and Seljak (2015); Raccanelli et al. (2010). This velocity term is usually ignored for mocks or surveys with small angular coverage Beutler et al. (2019); Beutler and McDonald (2021); Castorina et al. (2019), but it must be included in the full-sky limit Raccanelli et al. (2010); Benabou et al. (2024). The angular kernel corresponding to this velocity term is
@@
where $\alpha(x)\equiv\partial[\ln(x^{2}R(x))]/\partial\ln x$, and it becomes $\alpha=2$ for a uniform radial window in the redshift bin Yoo and Seljak (2015); Castorina and White (2018).
The mocks in Ref. Benabou et al. (2024) are generated with $f=0.71$ and $b_{1}=1.455$, and there is no redshift evolution within the redshift bin. All calculations assume a best-fit Planck 2018 cosmology Planck Collaboration VI (2020). The results for the mean of the multipole measurements from the 10,000 mocks with $L$ from 0 to 4 are shown in Fig. 1. We see that our theory predictions agree exceptionally well with the measured multipoles at the largest scales. The disagreements at the small scales ($k\gtrsim 0.06\,h/{\rm Mpc}$) are due to the voxel windows used in the simulations Benabou et al. (2024) which we do not include in our modeling.
IV Gaussian Covariance
We next express the Gaussian covariance of PSM in terms of the gSFB power spectrum. Compared to Ref. Wadekar and Scoccimarro (2020), which had the state-of-art results for analytical Gaussian covariance of the PSM, we will calculate the exact covariance without resorting to any plane-parallel or LOS approximation. The results in Ref. Wadekar and Scoccimarro (2020) are only applicable in the flat sky limit and on scales where the window function is subdominant, while our expression is exact and can be applied at the largest scales in particular for measuring PNG and GR effects. Using the Yamamoto estimator of Eq. 12, the Gaussian covariance reads Wadekar and Scoccimarro (2020):
@@ IV Gaussian Covariance
Using Wick contractions on the 4-point function and expressing $F_{L}(k)$ with Eq. 13, we can then follow the same procedure as the PSM case. A more detailed derivation will be given in Appendix B. To simplify the final expression, we further assume the separability of the window function and obtain:
$\textbf{C}^{\textup{G}}_{L_{1}L_{2}}(k_{1},k_{2})=(4\pi)^{4}\frac{(2L_{1}+1)(2L_{2}+1)}{I_{22}^{2}}\sum_{a,b,c,d,\ell_{1},\ell_{2}}i^{-a-c+b+d}(2a+1)\begin{pmatrix}a&L_{1}&b\\ 0&0&0\end{pmatrix}^{2}\begin{pmatrix}c&L_{2}&d\\ 0&0&0\end{pmatrix}^{2}$
$\qquad\Bigg{[}(2c+1)S_{b\ell_{1}d\ell_{2}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(-1)^{c+d}}(2d+1)S_{b\ell_{1}c\ell_{2}}\Bigg{]}C_{\ell_{1}}^{ad,{\rm R}}(k_{1},k_{2})C_{\ell_{2}}^{bc,{\rm R}}(k_{1},k_{2})\,,$ (16)
@@
Under the same set-up as our log-normal mocks with full-sky window, we plot an example of the analytical Gaussian covariance matrix using Eqs. 8 and 19 in Fig. 2, which shows prominent off-diagonal components with $k\lessapprox 0.02h/{\rm Mpc}$.